### v1

# Proprioceptive Image: An Image Representation of Proprioceptive Data from Quadruped Robots for Contact Estimation Learning

Abstract
This paper presents a novel approach for representing proprioceptive time-series data from quadruped robots as structured two-dimensional images, enabling the use of convolutional neural networks for learning locomotion-related tasks. The proposed method encodes temporal dynamics from multiple proprioceptive signals, such as joint positions, IMU readings, and foot velocities, while preserving the robot’s morphological structure in the spatial arrangement of the image. This transformation captures inter-signal correlations and gait-dependent patterns, providing a richer feature space than direct time-series processing. We apply this concept in the problem of contact estimation, a key capability for stable and adaptive locomotion on diverse terrains. Experimental evaluations on both real-world datasets and simulated environments show that our image-based representation consistently enhances prediction accuracy and generalization over conventional sequence-based models, underscoring the potential of cross-modal encoding strategies for robotic state learning. Our method achieves superior performance on the contact dataset, improving contact state accuracy from 87.7% to 94.5% over the recently proposed MI-HGNN method, using a 15 times shorter window size.
I INTRODUCTION
Deep learning has achieved remarkable success in domains where sequential or high-dimensional data can be transformed into visual representations suitable to convolutional architectures. In speech recognition, for instance, raw audio signals are often converted into spectrograms, two-dimensional time–frequency images, that serve as powerful inputs to convolutional neural networks (CNNs) and other image-based models. These approaches have yielded state-of-the-art performance across a range of tasks, from keyword spotting to end-to-end transcription [1] [2] [3]. Similarly, in cybersecurity, software files and their execution behavior have been rendered as grayscale or RGB images, where visual patterns detectable by CNNs facilitate malware detection [4]. These data-to-image transformations have unlocked new pathways by leveraging the spatial feature extraction strengths inherent to image-oriented deep learning.
Inspired by these cross-domain successes, this work investigates the transformation of proprioceptive time-series data collected from multiple proprioceptive signals of a quadruped robot into two-dimensional images that encode both temporal dynamics and robot morphology. By arranging proprioceptive signal readings according to the physical layout of the robot and incorporating time as an additional spatial dimension, the resulting images preserve inter-signal correlations and gait-dependent patterns in a form optimized for CNN processing. This approach provides a richer and more spatially coherent feature space than models trained directly on raw time-series inputs, enabling more effective learning of complex locomotion tasks. We demonstrate its potential in the context of contact estimation, where accurate prediction of foot–ground interactions is essential for stable and adaptive quadruped locomotion [5] [6].
Figure 1 illustrates three time steps of a quadruped robot walking over uneven terrain. For each time step, the corresponding proprioceptive images are shown, derived from base angular velocity from IMU, joint position from encoders, and estimated foot position proprioceptive signals. These images highlight how the representation varies across different environments, providing a compact encoding that supports learning of fundamental tasks such as foot contact estimation.
Figure 1: Examples of proprioceptive images over three time steps of a quadruped robot walking on uneven terrain. Each image encodes angular velocity, joint position, and foot position, illustrating how the representation varies across environments and supports learning of tasks such as foot contact estimation.
I-A Related Work - 1D time series to image conversion
To apply computer vision methods to time-series analysis, Wang and Oates [7] proposed the Gramian Angular Field (GAF) and Markov Transition Field (MTF) encodings, which enable CNN-based classification. GAF constructs a Gramian matrix [8] where elements capture trigonometric relations between time intervals, offering a bijective mapping that preserves temporal correlations and allows approximate reconstruction of the original series.
In contrast, the MTF encodes the dynamic transition statistics of a time series by discretizing it into quantile bins and constructing a transition matrix over time [7]. While MTF effectively represents multi-span transition probabilities, including self-transitions along the diagonal, its performance is sensitive to the choice of bins, complicating its use with heterogeneous signal modalities. Moreover, both MTF and GAF suffer from rate invariance, producing identical images for increasing and decreasing sequences with the same rate.
The Recurrence Plot (RP) [9] captures temporal correlations by marking state recurrences, enabling the diagnosis of drift, periodicity, or chaotic dynamics. However, it is less sensitive to subtle variations, which limits its effectiveness.
Transforming time series into images such as GAF, MTF, or RP enables CNN-based spatial feature learning and has shown strong performance across benchmarks [10, 11, 12]. Multimodal fusion through meta-feature fusion, attention, or concatenation further improves accuracy by capturing complementary static, dynamic, and recurrence patterns, with successful applications in manufacturing defect detection and healthcare heartbeat classification [13, 14, 15]. Overall, fused image encodings combined with tailored CNN/attention architectures remain a practical and effective strategy for time-series classification [16, 17].
In this work, we adapt the motivation of prior studies on converting 1D time-series data into 2D image representations for feature extraction via CNNs to the context of quadruped robotics. We propose a novel encoding technique that leverages proprioceptive signal data and the robot’s morphology to generate structured image representations, enabling the robot to learn tasks such as foot contact estimation.
I-B Related Work - Contact Estimation
Contact estimation for quadruped robots is a fundamental perception problem, as contact measurements are not always available [6]. However, it is a hard problem due to sensor noise, terrain variability, and other factors such as slippage.
A common method to estimate contact is to use a threshold value with the estimated or measured ground reaction forces (GRF) [18]. However, this approach often fail to generalize across robots or environments. To address these challenges, diverse approaches leverage probabilistic modeling, sensor fusion, and system dynamics. Camurri et al. [5] introduce a probabilistic method that estimates ground contact by using the GRF computed from joint positions, velocities, and efforts. Other frameworks, such as the Interacting Multiple-Model Kalman Filter (IMM-KF) [19], model the robot’s movement as a switched system where each mode represents a different set of feet in contact. Alternatively, Maravgakis et al. [20] propose a probabilistic technique that uses Kernel Density Estimation (KDE) over batches of IMU data to compute the probability of stable six degree-of-freedom contact. Other methods include the Dual $\beta$-Kalman Filter [21], which confronts the challenges of foot slippage and variable leg length.
Traditional model-based and thresholding approaches to contact estimation remain sensitive to noise and require empirical tuning, limiting robustness on lightweight robots or varying terrains. To overcome these challenges, recent work has moved toward data-driven techniques. For instance, Youm et al. [22] use a Neural Measurement Network to output contact probability, enhancing IEKF performance under slip conditions. Similarly, Lee et al. [23] estimate contact probabilistically from dynamics and joint torques, with a neural compensator correcting residual errors. Beyond these hybrid strategies, fully learning-based methods have also emerged. For instance, Lin et al. [24] proposed Contact-CNN, a multi-modal deep model that relies solely on proprioceptive signals (joint encoders, IMU, kinematics) to classify foot–ground contact across diverse gaits and terrains, avoiding explicit contact sensing altogether.
Ordonez-Apraez et al. [25] introduced Discrete Morphological Symmetries (DMSs), a group-theoretic symmetry groups of dynamical systems. Building on this, they proposed ECNN, a variant of Contact-CNN that applies sparse equivariant convolutions and batched augmentation to enforce hard equivariance, making large architectures tractable and improving leg-level F1 and test loss. Also, MI-HGNN [26] introduces a morphology-informed heterogeneous graph neural network that constructs a graph from a robot’s kinematic structure (nodes being the base/joints/feet, and edges representing links) to constrain learning for contact perception problems. Empirically, MI-HGNN attains substantially higher contact-detection accuracy (mean 87.7% vs 78.8% from ECNN [25]) on the Contact Data Set [24].
Although recent data-driven methods [24, 25, 26] have achieved promising results, they still rely on raw proprioceptive data streams as input. In contrast, our approach introduces a novel representation to create unique image patterns to describe quadruped robot sensing. To the best of our knowledge, this is the first method in robotics that converts proprioceptive signals into image representation.
I-C Contributions
The contributions of the paper can be summarized as follows:
  • A novel representation of proprioceptive signals into 2D images, with the name Proprioceptive image (PI). The core objective of the PI representation is to encode both temporal dynamics and morphological characteristics of quadruped robots. This image representation can be explored in learned-based approaches to perform complex tasks in robotics.
  • We propose a CNN architecture that is able to learn PI information and perform contact estimation for quadruped robots.
  • We show the results of the PI trained models in real-world datasets and in simulated environments, and compared with state-of-the-art methods. In the contact dataset from [24], our method improved contact state accuracy from 87.7% to 94.5% over the recently proposed MI-HGNN method [26], using a 15 times shorter window size.
II METHODOLOGY
In this section, we present the four key components of the PI formulation and explain how they are fused to construct the final PI representation, which is tailored to the robot’s morphology. Each component captures essential temporal and statistical characteristics of a single proprioceptive state and encodes them into a sub-image through well-defined mathematical transformations. Specifically, it is extracted and encoded into image form the following elements from the signal data stream: the slope dynamics in Section II-B, spike patterns in Section II-C, modified GAF polar distance in Section II-D, and the global-aware local shift in Section II-E. Figure 2 shows how to generate the Proprioceptive Image from proprioceptive signal data.
II-A Preliminaries
Input Values: Let $s=\{s_{i}\}^{T}_{i=1}$ be a discrete 1D time series sampled from a proprioceptive signal (estimated or measured by a sensor). At each timestamp $t$, we define a sliding window with length size $w$ as $x_{t}=[s_{t-w+1},\dots,s_{t}]\in\mathbb{R}^{w}$.
Figure 2: Pipeline for constructing a Proprioceptive Image. Proprioceptive signal streams are grouped in triplets (X, Y, Z for trunk signals; HAA, HFE, KFE for leg signals) to form three-channel RGB inputs. Each windowed signal is processed by a temporal feature encoder that generates image-based encodings of slope dynamics, spike patterns, and GAF. In parallel, the local range of each signal triplet is used to create the final PI sub-image, named Global-Aware Local Shift Image, which is spatially concatenated with the PI Temporal Feature Encoder into a square image. Leg PIs are further arranged in a $\begin{bmatrix}LF&RF\\ LH&RH\end{bmatrix}$ layout to represent all four legs, while trunk PIs are used individually. Yellow circles labeled ‘C’ indicate channel concatenation, and circles labeled ‘S’ denote spatial concatenation.
Global Bounds: The global bounds are defined as the maximum and minimum values of the proprioceptive signals outputs $\mathcal{G}=[s^{global}_{min},s^{global}_{max}]$ (usually obtained from the sensor datasheet). $\mu_{\mathcal{G}}$ is defined as the average of the global bounds $\mu_{\mathcal{G}}=(s^{global}_{max}+s^{global}_{min})/2$. The normalized sliding window sequence, computed with respect to the global bounds, is defined as $x_{t}^{\mathcal{G}}=[s_{t-w+1}^{\mathcal{G}},\dots,s_{t}^{\mathcal{G}}]\in\mathbb{R}^{w}$, where each $s_{i}^{\mathcal{G}}\in[-1,1]$ represents the globally normalized signal value at time $i$ computed as:
$s_{i}^{\mathcal{G}}=\frac{(s_{i}-s^{global}_{max})+(s_{i}-s^{global}_{min})}{(s^{global}_{max}-s^{global}_{min})}$ (1)
Local Range: For each proprioceptive signal of interest, it is computed a pair of normalized deviation measures $\delta$. These values quantify how far the current local operating range $\mathcal{Y}=[s^{local}_{min},s^{local}_{max}]$ of the signal is shifted relative to its global reference range. The local range is estimated from a sliding window by taking the lower and upper percentiles (10th and 90th) and expanding them by a small margin to ensure robustness to noise. Given the global average $\mu_{\mathcal{G}}$ and global bounds $\mathcal{G}$, the normalized deviation $\delta=[\delta_{min},\delta_{max}]$ is defined as:
[…]
Quadruped Robot Morphology: The quadruped robot consists of a rigid trunk and four identical legs — left-front (LF), right-front (RF), left-hind (LH), and right-hind (RH). Each leg has three actuated joints: hip abduction/adduction (HAA), hip flexion/extension (HFE), and knee flexion/extension (KFE). Each joint provides proprioceptive feedback, including torque (or torque estimates from motor currents), angular position, and velocity. The trunk houses an inertial measurement unit (IMU) for measuring accelerations, angular velocities, and estimating body orientation.
PI Concatenation: The primary objective of each PI is to capture key dynamics of a data stream and encode them into a compact image representation. To also incorporate the robot’s morphology, individual PIs are concatenated both channel-wise and spatially to form a final representation for a given proprioceptive signal type (e.g., torque, joint position, joint velocity). PIs are grouped into leg PIs and trunk PIs. Leg signals include per-leg proprioceptive signals such as joint positions, joint velocities, foot positions, foot velocities, and GRFs. Trunk signals, such as the data from IMUs, provide measurements of the trunk’s motion in space (angular velocity and linear acceleration), which serve as global indicators of the robot’s state. For example, the torque PI for a single leg $\textsc{PI}^{Leg}\in\mathbb{R}^{2w\times 2w\times 3}$ is formed by concatenating channel-wise the PIs of its three joints (or three linear dimensions in the case of foot positions and foot velocities):
$\textsc{PI}^{Leg}_{Torque}=\begin{bmatrix}(\textsc{PI}^{\text{HAA}}_{Torque},\textsc{PI}^{\text{HFE}}_{Torque},\textsc{PI}^{\text{KFE}}_{Torque})\end{bmatrix}$ (3)
and the full-body torque PI is obtained by arranging the per-leg torque PIs spatially:
$\textsc{PI}_{\text{Torque}}=\begin{bmatrix}\textsc{PI}^{LF}_{Torque}&\textsc{PI}^{RF}_{Torque}\\ \textsc{PI}^{LH}_{Torque}&\textsc{PI}^{RH}_{Torque}\end{bmatrix}\in\mathbb{R}^{4w\times 4w\times 3}$ (4)
Trunk PIs are constructed by concatenating the components of each dimension channel-wise. The concatenation technique is used in all key components of the PI except the Global-Aware Local Shift.
II-B Slope Dynamics
Given the temporal window of scalar values $x_{t}$, we encode three dynamical descriptors — mean slope, jerk, and peak density — into a spatial Gaussian pattern modulated by concentric ripples. More precisely, we fit a linear model $x_{i}=\beta_{t}i+b$ over the window $w$ to compute the slope $\beta_{t}\in\mathbb{R}$ using least-squares and clipped to be [1,1][-1,1]:
$\hat{\beta}_{t}=\text{clip}\Big(\frac{\beta_{t}}{\sigma(x)+\epsilon},-1,1\Big)$ (5)
where $\sigma(\cdot)$ is the standard deviation of the window, and $\epsilon$ ensures numerical stability. The normalized slope $\hat{\beta}$ determines the horizontal displacement of the blob. The jerk is computed as the absolute value of the second time-derivative,
$\mathcal{J}=\frac{1}{w-2}\sum_{t=0}^{w-3}|(x_{t+2}-x_{t+1})-(x_{t+1}-x_{t})|$ (6)
and normalized as $\hat{\mathcal{J}}=\text{clip}(\frac{\mathcal{J}}{\sigma(\dot{x}_{t})+\epsilon},0,1)$ by the velocity deviation $\sigma(\dot{x})$, which defines the vertical displacement of the blob. The peak density is defined as the ratio between the number of local maxima and the window size, scaled to $[0,1]$:
$\text{peak}=\text{clip}\Big(\frac{\#peaks}{w}\cdot 5,0,1\Big)$ (7)
which controls the ripple frequency $f_{ripple}=2+\lfloor 6\cdot\text{peak}\rfloor$. These normalized features are mapped to the center of a Gaussian blob, whose coordinates $c_{x}$ and $c_{y}$ are given by:
[…]
Let $R_{i,j}=\sqrt{(i-c_{x})^{2}+(j-c_{y})^{2}}$ be the radial distance from the blob center, the base Gaussian is computed as:
$G_{i,j}=exp{\Bigg(-\frac{R_{i,j}^{2}}{2\sigma_{g}^{2}}\Bigg)}$ (9)
with Gaussian spread of $\sigma_{g}=w/6.25$. To introduce the ripple texture, we apply a cosine modulation:
$\rho_{i,j}=0.5+0.5\cos\Bigg(\frac{f_{ripple}R_{i,j}}{\sigma_{g}}\Bigg)$ (10)
and form the final pattern $A=G\odot\rho$, normalized to the interval $[0,1]$ and scaled to 8-bit intensity. In this representation, the horizontal position of the blob encodes trend direction (slope), the vertical position encodes motion irregularity (jerk), and the ripple density reflects oscillatory content (peak density), providing a compact visual signature of local signal dynamics. Figure 3 presents three example images generated from different input signals.
Figure 3: Examples of Slope Dynamics images from simulated joint position signal of a quadruped robot trotting.
II-C Spikes Patterns
Given a temporal window $x_{t}$ from the proprioceptive sensor, we form the pairwise difference matrix $\Delta_{i,j}=x_{j}-x_{i}\in\mathbb{R}^{w\times w}$ and suppress noise via an adaptive gating threshold based on robust variability. Specifically, we compute the median $m_{x}=\text{median}(x)$ and the median absolute deviation (MAD)
$MAD_{x}=\text{median}(|x_{t}-m_{x}|)+\epsilon$ (11)
and set the gating threshold as $th_{\text{MAD}}=\alpha\text{MAD}_{x}$, where $\alpha>0$ controls the sensitivity (smaller values detect more subtle changes, and larger values restrict detection to stronger transitions). A binary gate
$\mathcal{M}_{i,j}=\mathbb{I}(|\Delta_{i,j}|>th_{MAD_{x}})$ (12)
selects only pairs exceeding this adaptive threshold. For the masked upper-triangular entries $(i<j,\mathcal{M}_{i,j}=1)$, we compute their median $m_{\Delta}$ and $\text{MAD}_{\Delta}$ to form the following robust score
$Z_{i,j}=\frac{\Delta_{ij}-m_{\Delta}}{MAD_{\Delta}}$ (13)
Scores are symmetrically clipped to [-3,3], following the three-sigma principle [27] to limit extreme outliers while maintaining contrast for typical values. The final spike image $I\in[0,255]^{w\times w}$ is obtained as
$C_{i,j}=\begin{cases}128+127\frac{\text{clip}(Z_{ij},-3,3)}{3}&\mathcal{M}_{ij}=1\\ 128&\mathcal{M}_{ij}=0\end{cases}$ (14)
where bright pixels $(\approx 255)$ represent strong positive spikes, dark pixels $(\approx 0)$ strong negative spikes, and mid-gray $(\approx 128)$ either neutral changes or subthreshold differences. This adaptive MAD-based construction yields a symmetric, noise-resilient spike matrix that adjusts to local variability while preserving sensitivity to significant temporal transitions. Figure 4 illustrates three examples of proprioceptive images obtained from distinct input sequences.
Figure 4: Examples of Spikes pattern images from simulated joint position of a quadruped robot trotting.
II-D Polar Distance
In this sub-image, we employ a modified version of the GAF representation. While the standard GAF inherently captures static temporal information through its unique polar encoding, it suffers from ambiguity in distinguishing between increasing and decreasing trends. To overcome this limitation, we replace the values along the main diagonal of the GAF matrix with the globally normalized sliding window sequence $x_{t}^{\mathcal{G}}$, subsequently scaled to the pixel intensity range [0, 255]. This substitution embeds directional information directly into the matrix diagonal values, enhancing the interpretability of the encoded temporal dynamics. Figure 5 shows three representative proprioceptive images corresponding to different input conditions. Eq. 15 describes the GAF function:
$D=GAF(x_{t},x_{t}^{\mathcal{G}})\in\mathbb{R}^{w\times w}$ (15)
Figure 5: Example of GAF images from simulated joint position signal of a quadruped robot trotting.
II-E Global-Aware Local Shift
We propose the cymatic encoding, inspired by physical vibration patterns observed on resonating plates and membranes, where standing wave interference produces stable, symmetric geometric patterns that depend on the excitation frequencies and phases. Such patterns are a natural 2D analog of harmonic decomposition in 1D signals: radial and angular wave modes act like spatial Fourier components, while their interference encodes complex multi-frequency relationships in a compact visual form.
This sub-image is the only one that does not follow the PI concatenation method as the other three sub-images. Instead of inputting the local temporal window and concatenating the PI images for each joint/axis, we input the normalized deviations $\delta$ of three joint/axis proprioceptive signals as a 6-dimensional descriptor vector that controls the spatial pattern parameters. The first four components set the radial frequency $k_{r}\in(1,4)$, angular frequency $k_{t}\in(1,4)$, and their respective phases $\phi_{r}$$\phi_{t}\in[0,2\pi)$. These determine the number of concentric rings and azimuthal lobes, directly modulating spatial symmetry. The fifth component sets a modulation scale $\alpha\in[0,1]$, controlling the amplitude of a secondary interference field, while the sixth is a blend weight, $m\in[0,1]$ determining how strongly the primary and secondary fields combine.
The 2D spatial domain $(X,Y)\in[-1,1]^{2}$ is converted to polar coordinates
[…]
and two wave-interference components are defined as $\mathcal{C}_{1}=\sin(k_{r}\mathcal{R}+\phi_{r})\cos(k_{t}T+\phi_{t})$ and $\mathcal{C}_{2}=\sin(k_{r}\mathcal{R}-k_{t}T)$.
The final cymatic field is a convex combination given as
$\mathcal{C}=(1-m)\mathcal{C}_{1}+m\alpha\mathcal{C}_{2}$ (17)
masked outside the unit disk and normalized to [0, 255] for image representation. This mapping is special since it translates a low-dimensional descriptor into a high-complexity, structured, and interpretable 2D pattern, whose symmetry, periodicity, and interference texture vary smoothly with its parameters. Unlike arbitrary texture generators, cymatic patterns are rooted in physical wave phenomena, yielding repeatable geometric signatures that are both parameter-sensitive and visually distinct. This makes them particularly well-suited for encoding multi-feature signal descriptors into images for downstream convolutional processing, as small variations in the descriptor lead to smooth but discriminative changes in the spatial pattern. Figure 6 depicts three sample images illustrating how varying inputs produce distinct representations.
Figure 6: Examples of Global-Aware Local Shift images from simulated joint position signal of a quadruped robot trotting.
II-F Contact Estimation
Following [24], contact estimation is framed as a multi-class classification problem, where each binary foot–ground configuration corresponds to a discrete class. To address this, we propose ConcatCNN, a multi-image convolutional architecture that jointly encodes and fuses proprioceptive image representations. Each input image is processed by a shared CNN encoder with two convolutional layers (kernel size $3\times 3$, padding 1), each followed by Group Normalization [28] and Leaky ReLU activation [29], with a $2\times 2$ max-pooling operation for downsampling. Leaky ReLU mitigates the risk of dead neurons in sparse proprioceptive inputs, while ReLU activations are retained in the fully connected layers for stability and efficiency. Group Normalization was chosen over Batch Normalization for its robustness to small batch sizes and heterogeneous signals streams, ensuring consistent feature scaling. The shared encoder outputs feature maps of size $16\times 5\times 5$ per image, which are concatenated into a fused representation of size $16N\times 5\times 5$, where $N$ denotes the number of proprioceptive images (PIs) provided as input. This representation is refined by a convolutional fusion layer (kernel size $3\times 3$, 32 channels), flattened, and passed through a two-layer fully connected network (hidden dimension 128, ReLU activation) with dropout [30] for regularization. The final output layer predicts the contact state across the predefined classes. Overall, ConcatCNN enables efficient joint feature extraction and fusion by leveraging parameter sharing in the encoder and channel-wise concatenation for cross-signal integration, thereby enhancing the robustness of contact state classification. The complete architecture is depicted in Fig. 7, which highlights the shared encoders, fusion stage, and final prediction layers.
Figure 7: Architecture of ConcatCNN for quadruped contact estimation. Each proprioceptive image is encoded by a shared CNN, and the resulting features are stacked and fused to jointly learn from multiple proprioceptive signals. In the case of fusing trunk and leg PIs, the concatenated four legs PI are separated. The fused representation is flattened and mapped through fully connected layers to predict discrete foot–ground contact states.
III RESULTS
III-A Real-World Dataset Results
In [24], a comprehensive open-source contact dataset was developed using the MIT Mini Cheetah robot, comprising approximately one million labeled data points collected over eight different terrains (asphalt, concrete, forest, grass, middle pebble, small pebble, rock road, and sidewalk) and multiple gaits. This dataset incorporates both contact and non-contact scenarios to enhance generalization. Ground truth contact labels were generated with a self-supervised algorithm that analyzes filtered foot height signals, thereby circumventing the need for additional external sensors during data collection.
The dataset was converted into PI representations using a sliding window size of $w=10$, generating images $I\in\mathbb{R}^{20\times 20\times 3}$. The proposed contact estimation model, described in Section II-F, was trained with the same proprioceptive signal modalities as [24] to ensure a fair comparison. These modalities include joint positions, joint velocities, foot positions, foot velocities, trunk angular velocity, and trunk linear acceleration. Performance was assessed in terms of overall accuracy, F1 score, per-leg accuracy, and LegAvg accuracy (the mean of the individual leg accuracies). Table I summarizes the results across the different test sets, comparing our PI-based model against the approach in [24]. The PI model achieved consistently higher overall accuracy and F1 score, with the latter balancing precision and recall to account for potential class imbalance, while per-leg and LegAvg accuracies were nearly identical across methods, differing only marginally. This indicates that PI representations yield more robust global classification performance while preserving per-leg accuracy at a high level, despite relying on sequences of only 10 time steps compared to 150 in [24].
Dataset Method Acc(%) F1 LF(%) RH(%) RF(%) LH(%) Leg Avg(%) Concrete PI CNN 94.05 0.9337 98.19 97.33 98.03 98.32 97.97 Lin et al. [24] 93.04 0.4145 97.42 97.12 97.49 97.58 97.40 Forest PI CNN 91.09 0.8980 96.49 97.66 96.89 97.07 97.03 Lin et al. [24] 89.95 0.5416 96.39 97.54 97.11 96.81 96.96 Grass PI CNN 94.27 0.9333 98.21 98.08 98.11 98.17 98.14 Lin et al. [24] 92.87 0.5348 97.58 97.31 97.85 97.45 97.55 Asphalt Road PI CNN 95.20 0.9520 98.63 98.42 98.71 98.88 98.66 Lin et al. [24] 94.13 0.6023 97.74 98.19 98.46 98.28 98.17 Middle Pebble PI CNN 93.18 0.9270 97.77 97.49 97.54 97.70 97.62 Lin et al. [24] 92.45 0.4861 97.84 96.90 97.16 97.94 97.46 Small Pebble PI CNN 91.80 0.9032 97.82 96.11 97.94 97.62 97.37 Lin et al. [24] 90.65 0.4953 96.88 96.41 97.85 97.27 97.10 Sidewalk PI CNN 95.21 0.9446 98.18 97.61 98.20 98.49 98.12 Lin et al. [24] 93.71 0.4174 97.43 97.68 97.65 97.06 97.53 Rock Road PI CNN 93.72 0.9313 97.69 98.57 97.39 98.79 98.11 Lin et al. [24] 92.32 0.5338 97.32 97.67 97.29 97.55 97.46
TABLE I: Contact Estimation comparison across datasets from [24].
We also evaluated our approach using the same metrics as [25] [26], namely the foot-wise binary F1-score, the averaged F1-score (mean of the four leg-specific F1-scores), and the contact state accuracy over 16 possible states. The latter is a stricter metric than averaged accuracy, as a prediction is counted as correct only when the contact state of all four feet is simultaneously correctly predicted. To compute this metric for our model, the foot-wise predictions were converted into a 16-state representation. In addition, the real-world dataset provided by [24] was transformed into PIs using a sliding window of size $w=10$, consistent with the results reported in Tab. I. Unlike the original setup, however, our model was trained on the combined training and validation sets and evaluated on the combined test sets. The experimental results, presented in Fig. 8, demonstrate that our method outperforms all state-of-the-art data-driven methods (MI-HGNN [26], ECNN [25], CNN-aug [25], and CNN [24]) in the Legs-Avg F1 (from 0.9356 of MI-HGNN to 0.9732) and overall model accuracy (from 87.7% of MI-HGNN to 94.5%) with smaller standard deviation. These results show the potential of this representation by obtaining better accuracy with a less complex NN architecture with relation to other methods [25] [26].
Figure 8: Contact detection results on the real-world Mini- Cheetah contact dataset [24]: classification performance of five models on the unseen test set, trained with the entire training set. The mean and standard deviation across 8 runs are reported and compared with MI-HGNN [26], ECNN [25], CNN-aug [25] and CNN [24].
III-B Simulation Results
To further evaluate our PI formulation, we generated a dataset in the MuJoCo simulator [31] using a 60kg Unitree B2 quadruped robot. The dataset includes three environments (flat, uneven, and pyramid), two gaits (trot and crawl), and both stable ($\mu=1.0$) and slippery ($\mu=0.2$) terrain conditions. The uneven and pyramid environments were randomized in relation to the boxes positions and the height of the pyramid. The robot traversed these environments under a gradient-based MPC controller developed in [32], with trunk linear velocities ranging from –1 to 1 m/s and angular velocities from –0.3 to 0.3 rad/s. Proprioceptive signals were recorded, including joint positions, joint velocities, foot positions, foot velocities, IMU signals and GRFs with Gaussian noise added to better approximate real-world sensing conditions. Contact estimation labels were directly extracted from the simulator and used as ground truth for training and evaluation. In total, the dataset comprises more than 1.5M synced data, organized by robot gait and environment stability. Each subset was further divided into training, validation, and test splits to enable fair comparison between our PI formulation, the approach of [24], and a baseline method based on GRF threshold [6, 18]. Figure 9 illustrates examples of scenes from the simulated dataset across the different environments.
(a) Uneven Terrain
(b) Flat Terrain
(c) Stairs Terrain
Figure 9: Mujoco Simulated Environment.
Test Dataset Method Trained Acc(%) F1 LF(%) RH (%) RF (%) LH (%) Leg Avg (%) Trot - Stable PI CNN Stable 99.46 0.9600 99.86 99.86 99.90 99.80 98.85 PI CNN Fused 99.39 0.9565 99.80 99.83 99.88 99.78 99.82 Lin et al. [24] Stable 96.69 0.5840 99.00 98.96 99.22 99.14 99.08 Lin et al. [24] Fused 96.79 0.7291 98.99 98.91 98.88 99.05 98.96 GRF TH - - - 69.741 69.58 69.41 69.62 69.95 Trot - Slippery PI CNN Stable 88.41 0.7165 96.36 96.09 95.99 95.50 95.98 PI CNN Fused 96.16 0.8925 98.96 99.13 98.85 98.79 98.93 Lin et al. [24] Stable 86.69 0.3510 94.31 94.62 95.67 95.64 95.06 Lin et al. [24] Fused 91.24 0.5708 96.85 96.67 97.71 97.26 97.12 GRF TH - - - 66.14 65.75 64.81 65.22 65.485 Trot - Fused PI CNN Stable 93.77 0.7899 98.05 97.93 97.89 97.58 97.86 PI CNN Fused 97.73 0.9103 99.36 99.47 99.35 99.27 99.36 Lin et al. [24] Stable 91.45 0.3893 96.55 96.74 97.32 97.16 96.94 Lin et al. [24] Fused 93.61 0.5923 97.84 97.73 98.05 98.12 97.93 GRF TH - - - 67.93 67.66 67.10 67.105 67.53 Crawl - Stable PI CNN Stable 98.85 0.9982 99.76 99.71 99.67 99.70 99.71 PI CNN Fused 98.87 0.9462 99.73 99.73 99.68 99.71 99.71 Lin et al. [24] Stable 96.73 0.7347 99.03 99.03 99.28 99.19 99.13 Lin et al. [24] Fused 96.17 0.7595 99.01 98.66 99.20 99.11 99.00 GRF TH - - - 82.45 82.81 81.62 81.49 82.09 Crawl - Slippery PI CNN Stable 83.48 0.6308 95.64 96.15 94.39 94.07 95.06 PI CNN Fused 93.18 0.8575 98.62 98.36 98.11 97.57 98.16 Lin et al. [24] Stable 78.48 0.4098 92.89 92.14 94.94 93.78 93.44 Lin et al. [24] Fused 84.34 0.6021 94.75 95.10 96.40 96.21 95.61 GRF TH - - - 77.54 78.82 74.28 74.09 76.18 Crawl - Fused PI CNN Stable 94.26 0.7685 98.52 98.64 98.10 98.03 98.32 PI CNN Fused 97.17 0.8934 99.40 99.32 99.21 99.08 99.25 Lin et al. [24] Stable 90.24 0.5374 96.71 96.54 97.75 97.49 97.12 Lin et al. [24] Fused 92.79 0.6840 97.55 97.71 98.38 98.34 97.99 GRF TH - - - 79.96 80.87 78.18 77.71 79.18
TABLE II: Contact Estimation comparison across datasets from Mujoco. Stable Data is the environment without slippery terrain, and Fused Data is the combined datasets of slippery and stable terrains.
Table II presents a comparison among three approaches: our PI-based CNN, the method of Lin et al. [24], and a threshold-based GRF baseline. The MuJoCo dataset is used with the same window sizes as described in Section III-A. Both models were trained on the stable environment and evaluated on three test sets: pure stable, slippery, and a fused (stable + slippery) dataset. The contact threshold was obtained by sweeping candidate values over the dataset’s GRF in the Z axis and selecting the value that maximized agreement with the ground-truth contact labels. The PI-CNN consistently outperformed the baseline across all test conditions, including the slippery environment, despite never being exposed to slippery data during training. These results highlight that PI provides a richer and more discriminative feature representation, enabling superior learning compared to directly using raw proprioceptive signals.
III-C Implementation Details
All tests were performed on a computer with Intel i9 CPU with 64 GB of RAM and NVIDIA RTX3090 GPU running Ubuntu 20.04 LTS Linux. A dropout rate of 0.3 was applied in the fully connected layers as an additional form of regularization. Training was carried out using Cross-Entropy loss for multi-class classification, optimized with adamW [33], initial learning rate of $1\cdot 10^{-4}$, trained through 40 epochs with batch size of 64. To improve convergence, a ReduceLROnPlateau scheduler was employed, which decreased the learning rate by a factor of 0.2 when the validation loss failed to improve for 3 consecutive epochs.
IV Discussion
Although PI encodings provide rich learnable patterns, their creation is relatively slow (1.5 ms per image) due to the current Python implementation running on the CPU. We expect that a GPU-based implementation would substantially reduce generation time and make the formulation more practical for real-time robotic systems. The PI-based contact estimation model achieved a mean inference time of 8.5 ms on GPU, corresponding to a processing frequency of approximately 120 Hz.
V CONCLUSIONS
This paper presented a novel formulation of proprioceptive signals into an image representation. Each image representation encodes four important temporal information of dynamic measurement behavior and are concatenated based on the quadruped robot morphology. The PI formulation was evaluated on the contact estimation problem, using simple CNN architectures trained on both simulated and real-world data, achieving the best results in comparison with state-of-the-art methods. The PI is the first step toward deeper representations of robot sensing. For future work, we expect that the PI can be integrated into more advanced learning structures such as Graph Neural Networks similar to [26] and also applied PIs as input for Reinforcement Learning techniques. Moreover, combining a PI with exteroceptive sensors, such as cameras, could yield a unified representation of both the robot’s internal state and its external environment, further enhancing perception and state learning. Finally, we aim to explore the applicability of PI in other robotic platforms, including humanoids and manipulators.